OpenStudy (anonymous):

how would you solve this exponential equation (1/2)^1/30 =.9

OpenStudy (anonymous):

hmm this is not an equation to solve because there is no variable. maybe a typo?

OpenStudy (anonymous):

oh yeah sorry, its t/30 not 1/30

OpenStudy (anonymous):

ok we go right to the answer, ( just about) remember that to solve \[b^t=A\] for t write \[t=\frac{\ln(A)}{\ln(b)}\] so in this case you have \[\frac{t}{30}=\frac{\ln(.9)}{\ln(.5)}\]

OpenStudy (anonymous):

to get t you put \[t=\frac{30\ln(.9)}{\ln(.5)}\]

OpenStudy (anonymous):

then a calculator